Mass of potassium acid oxalate required to reduce 100ml of 0.1 Molar kmno4 is a g while to neutralise 100ml of 0.1 Molar NaOH is b g. What is a/b?

Grade:11

1 Answers

Vikas TU
6 years ago
For potassium acid oxalate,
(a/128)*2 = 100*0.1*5/1000
a => 128*5/200 => 3.2 g
For NaOH,
(b/40)*1 = 100*0.1*5/1000
b = 2 g
 
a/b = 3.2/2